a. F(S) = NI b. F(9) = sta c. F(S) = d. F(s) = zi,Find the Inverse Laplace Transform of the following functions:
Answer #1

Please give LIKE

foi?= (@) Given F(s) = 4 Inverse Laplace transform ie Ky FC)} = { 133 ast I Caplace transform tl 1) =4 By Formula 1419 = = La- FC9= 2 is St3 Poveste Laplace transforming of | 1f9} = 2 * 2+ (C) Given FC) = 2453 . $ 769} = {} ca le make let us - some... The Inverse Laplace transforms of f(s) [ {Polyt (d) Given FCG) = s S l ovence Laplace transforma effle} Leidensis] =fals